Samantha and Rachel had lunch at the mall. Samantha ordered three slices of pizza and two colas. Rachel ordered two slices of pizza and three colas. Samantha's bill was $6.00 and Rachel's bill was $5.25. What was the price of one slice of pizza? What was the price of one cola?
Answer by fractalier(6550)

Call the prices, p and c. Then we have
3p + 2c = 6
2p + 3c = 5.25
Now multiply the top by 2 and the bottom by 3 and subtract 'em...we get
6p + 4c = 12
-(6p + 9c = 15.75)
--------------------
-5c = -3.75
c = $.75
Then
3p + 2(.75) = 6
3p + 1.50 = 6
3p = 4.50
p = $1.50
$1.50 for pizza and $.75 for a coke...